Question :



Question :











 







Grade:12

1 Answers

Rahul verma
10 years ago
I have taken theta as @;Sum of roots,sin @ + cos @ = b/a ...........(1)product of roots,sin @ . cos @ = c/a............(2)Squaring (1),sin @ ^2  +  cos @ ^2  + 2.sin @.cos@ = b^2 / a^21 + 2c/a = b^2 / a^2a^2 -b^2 + 2ac = 0